On a ground trip of 540 km in each direction, Mr. Armstrong averaged 9 km/h more returning than he did going. If the return trip took 2 hours less than the trip going, find the time for each direction.

Answer by mananth(16949)

on hisway there speed x
on his return speed x + 9

Distance = same 540 km 1
original time – time with increased speed = 2 hour
t=d/t

540 / x - 540 / ( x + 9 ) = 2.00
LCD= x ( x + 9 )
multiply by LCD
540 ( x + 9 ) - 540 x = 2 x (
540 x + 4860 - 540 x = 2 X^2 + 18 x
4860 = 2 X^2 + 18 x
2 X^2 + 18 x- -4860 = 0
/2
x^2+9x-2430=0
x^2+54x-45x-2430=0
x(x+54)-45(x+54)=0
(x-45)(x+54)=0
x=45 OR -54
ignore negative
speed = 45 km/h
&
54 km/h